The length of a fenced rectangular garden is 6 feet longer then the width. If the length of the fenced used to surround the garden is 80 feet. What are the dimensions of the garden
Answer by josgarithmetic(39617)

First sentence, L=6+w.
Second sentence, 2w+2L=80.

The obvious substitution gives you 2w+2(6+w)=80; simplify and solve first for w.
Use the result of w to get the value for L.